What is 1/3(2 1/4 times 1/2 - 1/4) + 5/24?
Answer by rfer(16322)

1/3(9/4*1/2-1/4)+5/24
1/3(9/8-2/8)+5/24
1/3*7/8+5/24
7/24+5/24=12/24=1/2